The sides of a triangle all have integer lengths. two sides have lengths 7 and 10. if the largest and smallest possible perimeters for the triangle are L and S, the value of L +S is ???
Answer by Alan3354(69443)

Any side of a triangle has to be more than the difference of the other 2, and less than the sum of the other 2.
---------------
c > 10 - 7
c > 3
------
c < 10 + 7
c < 17
-------
The smallest integer value is 4, and the largest is 16.
S = 7 + 10 + 4 = 21
L = 7 + 10 + 16 = 33
-----------
L + S = 54